OpenStudy (anonymous):

How many moles of oxygen gas react to yield 0.200 mole water? C5H12(g) + 8 O2(g) = 5 CO2(g) + 6H2O(g)

OpenStudy (kaederfds):

\[\frac{ 8 }{ 6 } \times 0.2 = \frac{ 4 }{ 15 } moles\]

OpenStudy (kaederfds):

comparing the coefficient @tima82